6. Ordered: Kefzol 500 mg IM Q 6 H

Available: Kefzol 225 mg/ ml

Give:_________How many doses per day?

7. Ordered: Vistaril 100 mg IM Q 4 H PRN

Available: Vistaril 10ml vial containing 50 mg/ml

Give:_________

8. Ordered: Demerol 22mg IM Stat

Available: Demerol 25 mg/0.5 ml

Give:_________

9. Ordered: Hydrocortisone 8 mg IV Q 4 H

Available: Hydrocortisone 100 mg/2 ml

Give:_________

10. Ordered: Tobramycin 40 mg IM Q 8 h

Available: Tobramycin 60 mg/1.5 ml

Give:_________ How many doses per day? ___________
Answer #1

Ans) 6) Ordered: Kefzol 500 mg IM Q 6 H

Available: Kefzol 225 mg/ ml

225 mg = 1 ml

500 mg = x

x= 500/225 = 2.22 ml in 4 doses i.e 6th hrly

7) Ordered: Vistaril 100 mg IM Q 4 H PRN

Available: Vistaril 10ml vial containing 50 mg/ml

50 mg = 1 ml

100 mg = x

x= 100/50 = 2 ml

8) Ordered: Demerol 22mg IM Stat

Available: Demerol 25 mg/0.5 ml

25 mg = 0.5 ml

22 mg = x

x= 22 × 0.5/25 = 0.4 ml

9) Ordered: Hydrocortisone 8 mg IV Q 4 H

Available: Hydrocortisone 100 mg/2 ml

100 mg = 2ml

8 mg = x

x= 8 × 2/100 = 16/100 = 0.16 ml

10) Ordered: Tobramycin 40 mg IM Q 8 h

Available: Tobramycin 60 mg/1.5 ml

60 mg = 1.5 ml

40 mg = x

x= 40 × 1.5 /60 = 1 ml in 3 doses.